# Explicit expressions for the gamma vector leading to connections to upper/lower bounds and structural properties

### 20 Feb 2024 v1

@@ Abstract
We find an explicit formula for the gamma vector in terms of the input polynomial in a way that extends it to arbitrary polynomials. More specifically, we find explicit linear combination in terms of coefficients of the input polynomial (using Catalan numbers and binomial coefficients) and an expression involving the derivative of the input polynomial. In the case where the input is the $h$-polynomial of a simplicial complex, this gives an interpretation of the gamma vector as a measure of local-global contributions. We also apply them to connect signs/inequalities of (shifts of) the gamma vector to upper/lower bound conditions on coefficients of the input polynomial. Finally, we make use of the shape of the sums used to make these estimates and connections with intersection numbers to relate these properties of the gamma vector to algebraic structures.
@@ Introduction
The main object that we are concerned with is the gamma vector associated to (reciprocal) polynomial $h(t)$ of degree $n$, which is the unique polynomial $\gamma$ of degree $\lfloor\frac{\deg 2}{2}\rfloor$ such that
@@
Alternatively, it can be taken as the coefficients $\gamma_{i}$ used to write $h(t)$ as a linear combination of $t^{i}(1+t)^{n-2i}$. It was originally defined by Gal [16] in connection to real-rootedness questions of $h$-polynomials of flag spheres and had lower bound results with $h$-vectors in mind related to postivity of the gamma vector. Properties of this vector have since been involved in many different examples including those far beyond the original context [6].
Since the gamma vector of a polynomial is defined in terms of a functional equation involving the starting polynomial, it is natural to ask if there are explicit expressions in terms of coefficients of the original polynomial. We find explicit formulas for the gamma vector in terms of the coefficients of $h(t)$ in a way that confirms the connections to lower bounds and derivatives involving the input polynomial mentioned above. In addition, they can be used to extend the definition of the gamma vector to arbitrary polynomials. Concretely, the components of the gamma vector/coefficients of $\gamma(u)$ written as a linear combination of the coefficients of $h(t)$ have coefficients coming from products of Catalan numbers and binomial coefficients (Theorem 1.1).
@@
Continuing in this direction, we also consider interactions with algebraic structures. Since many of the examples we had considered earlier were connected to independence complexes and broken circuits of matroids (which can be express in terms of intersection numbers), we were motivated to consider an input polynomial with coefficients coming from intersections of divsiors on a variety. So, we considered the case where the input polynomial is the specialization $V(u)=(Au+B)^{d}$ of the volume polynomial associated to a pair of divsiors $A,B$ on a smooth projective variety. In general, the gamma vector has the same sign as $(-1)^{r}$ when the intersections $A^{k}B^{d-k}$ alternate in sign. Suppose that these intersections are nonnegative. When we are working with $\mathbb{P}^{n}$, we can relate the sign of the gamma vector to the relative “size” of $A$ compared to $B$ (since these correspond to integers) (3.3, 3.2). When the ratio is very large, the gamma vector is negative. If it is very small, it is alternating in sign like the negative ratio case. Positivity properties of divisors are also related to the bounds in a more general setting since the $A^{k}B^{d-k}$ form a log concave sequence when $A$ and $B$ are nef. This can be used to reduce upper and lower bound conditions related to alternating sums of positive multiples of such terms to those on the first or last pair of elements in a sequence. Note that the connections to upper/lower bounds studied in the previous section carry over to volume polynomials with some simplifications. Finally, we give an explanation of the $\mathbb{P}^{n}$ case in terms of Segre classes of vector bundles and make use of these classes to connect the rational functions giving components of the gamma vector in the derivative expression to positivity questions related to Segre classes (Example 3.6).
@@ 1.1 An explicit formula involving Catalan numbers, binomial coefficients, and derivatives
(Lagrange inversion formula, Theorem 5.4.2 on p. 38 of [25] and Theorem 2.1.1 on p. 214 of [17])
@@ 1.2 Local-global differences
Note that $r\leq\lfloor\frac{n}{2}\rfloor$. This means that $12r\leq n<n+$ 1 and $J^{\prime}(u)(u+1)^{2r}$ has a positive power of $u+1$ in the denominator and $(1+u)h^{\prime}(u)-nh(u)$ in the numerator. In the numerator, the coefficient of $v^{k}$ is $(k-1)h_{k-1}-(n-k)h_{k}$. Before starting more explicit computations, we make a heuristic comment on the interpreation of the derivative. In the case of $f$-polynomials, the derivative of the $f$-polynomial gives the sum of the $f$-polynomials of the links over the vertices of the given simplicial complex. This is a special case of the result below.
@@ 2.1 Bounds involving translations
$\frac{a_{k+2r-1}}{a_{k}}\leq\frac{k\binom{2r-k-1}{r-1}}{(k+2r-1)\binom{k+r-1}{r-1}}$
@@
Since $B(t)=C(t+1)$, we have that
@@
$\Longleftrightarrow\frac{a_{k+2r-1}}{a_{k}}$ $\leq\frac{k\binom{2r-k-1}{r-1}}{(k+2r-1)\binom{k+r-1}{r-1}}$
in order to have $\gamma_{r}\leq 0$ for even $r$ as well. This is a sort of upper bound condition. ∎
@@ 2.2 Upper/lower bounds and general gamma vectors
If $k\geq 1$, the coefficient of $b_{k}$ inside is the sum of the index $i=k-1$ and $i=k$ terms of the sum given by
@@
$=(-1)^{k-1}\left(k\left(\binom{d-r-k}{r-k}-\binom{d-r-k-1}{r-k-1}\right)+d\binom{d-r-k-1}{r-k-r}\right)$
@@
where we used the identity $\binom{p}{q}-\binom{p-1}{q-1}=\binom{p-1}{q}$ in the last line. The coefficient of $b_{0}$ inside the sum is $-d\binom{d-r-1}{r-1}$.
This means that
@@
      if $0\leq r\leq\frac{d}{3}$ or $r=\frac{d}{2}$. This also holds for $\frac{d}{3}<r<\frac{d}{2}$ if
@@
      for all $10\leq k\leq r-$ 1, then $\gamma_{r}\leq 0$.
@@
Some examples where Part 1 would apply is for include ±$\pm\overline{\chi}_{M}(t)$ and a translate ±$\pm\overline{\chi}_{M}(t+1)$ for reduced characteristic polynomials of matroids $M$ (with sign depending on the parity of the rank of $M$). Thinking about Part 2, note that there are As for Part 2b), we can think of face vectors where the entries increase quickly (Theorem 2.9). Examples for individual indices coming from independence complexes of matroids are in equality cases of upper bounds on face vectors of independence complexes of matroids in work of Castillo–Samper (Theorem 3.3 on p. 10 of [11]). We can move in between cases 2a) and 2b) by looking at reciprocal polynomials.
@@
Given the simplification above, it is clear that Part 1 holds. Also, Lemma 2.5 directly implies the statements in Part 2 for $r=\frac{d}{2}$ since $r\gamma_{r}$ is a positive multiple of the alternating sum of the $b_{i}$ in this case. For the remaining indices in Part 2, we start by noting that the binomial terms in each sum of Proposition 2.4 give a decreasing sequence as the index increases since
@@
Finally, we consider Part 2b). Note that the coefficients of the $b_{k}$ are decreasing in the second sum in Proposition 2.4. In order for the terms to be increasing, we need to have
@@
If $0\leq r\leq\frac{d}{3}$, the coefficients of the first sum of Proposition 2.4 are also decreasing. So, the condition for the terms of the first sum to increase is to have
@@ 2.3 Gamma vectors and (modified) $h$-vector representations
We will focus on the expression from Proposition 2.4 since want to look at sums of positive terms and the older formula may involve some negative terms. For the second sum in Proposition 2.4, it suffices to have $b_{k}=f_{k-1}(\Delta)$ for some simplicial complex $\Delta$. It remains to consider the first sum. The main issue is that have a sum $11\leq k\leq r-$ 1 (meaning that 1 or 2 terms are missing) and $Q=r$ instead of $r-1$. To “complete” this to something coming from a simplicial complex, we can either write $\binom{d-r-k-1}{r-k}=\frac{d-r-k-1}{r-k-1}\cdot\frac{d-2r-1}{r-k}$ and add one additional term for $k=0$ or keep the given binomial coefficient $\binom{d-r-k-1}{r-k}$ and add 2 additional terms for $k=0$ and $k=r$. We can combine these observations with the following result:
@@ 3 Intersection numbers and volume polynomials
Also, $r\gamma_{r}=h(P)$ for some auxiliary $(d-r-2)$-dimensional simplicial poset $P$ when $d$ is sufficiently large.
@@
The second statement is an application of Proposition 2.10. ∎
@@
    More general positivity properties are connected to alternating sums of multiples of positive intersections of the form $A^{k}B^{n-k}$ (e.g. in the context of gamma vectors). If $A$ and $B$ are nef, the sequence $a_{k}=A^{k}B^{d-k}$ forms a log concave sequence. Recall that a sequence $\{a_{k}\}_{k=1}^{N}$ is log concave if $a_{k}^{2}\geq a_{k-1}a_{k-1}$ for all $12\leq k\leq N-$ 1. Since this is equivalent to $\frac{a_{k}}{a}_{k-1}\geq\frac{a_{k+1}}{a_{k}}$, the upper bound condition $\frac{a_{i+1}}{a_{i}}\leq C$ holds for all $11\leq i\leq N-$ 1 if and only if $\frac{a_{1}}{a_{2}}\leq C$. Similarly, the lower bound condition $\frac{a_{i+1}}{a_{i}}\geq C$ holds for $11\leq i\leq N-$ 1 if and only if $\frac{a_{N}}{a_{N-1}}\geq C$.
@@
$r\gamma_{r}$`=′′$=^{\prime\prime}[u^{r-1}]\frac{dB^{d-1}(A-B)}{(1+u)^{d-2r+1}(1-\rho u)^{d-1}}.$

### 26 Feb 2024 v2

@@ Abstract
We find an explicit formula for the gamma vector in terms of the input polynomial in a way that extends it to arbitrary polynomials. More specifically, we find explicit linear combination in terms of coefficients of the input polynomial (using Catalan numbers and binomial coefficients) and an expression involving the derivative of the input polynomial. In the case where the input is the $h$-polynomial of a simplicial complex, this gives an interpretation of the gamma vector as a measure of differences in local and global contributions. We also apply them to connect signs/inequalities of (shifts of) the gamma vector to upper/lower bound conditions on coefficients of the input polynomial. Finally, we make use of the shape of the sums used to make these estimates and connections with intersection numbers to relate these properties of the gamma vector to algebraic structures.
@@ Introduction
The main object that we are concerned with is the gamma vector associated to (reciprocal) polynomial $h(t)$ of degree $n$, which is the unique polynomial $\gamma$ of degree at most $\lfloor\frac{\deg n}{2}\rfloor$ such that
@@
Alternatively, it can be taken as the coefficients $\gamma_{i}$ used to write $h(t)$ as a linear combination of $t^{i}(1+t)^{n-2i}$ when $h(t)$ is a reciprocal polynomial. It was originally defined by Gal [16] in connection to real-rootedness questions of $h$-polynomials of flag spheres and had lower bound results with $h$-vectors in mind related to postivity of the gamma vector. Properties of this vector have since been involved in many different examples including those far beyond the original context [6].
Since the gamma vector of a polynomial is defined in terms of a functional equation involving the starting polynomial, it is natural to ask if there are explicit expressions in terms of coefficients of the original polynomial. We find explicit formulas for the gamma vector in terms of the coefficients of $h(t)$ in a way that complements the connections to lower bounds and derivatives involving the input polynomial mentioned above. In addition, they can be used to extend the definition of the gamma vector to arbitrary polynomials. Concretely, the components of the gamma vector/coefficients of $\gamma(u)$ written as a linear combination of the coefficients of $h(t)$ have coefficients coming from products of Catalan numbers and binomial coefficients (Theorem 1.1).
@@
Continuing in this direction, we also consider interactions with algebraic structures. Since many of the examples we had considered earlier were connected to independence complexes and broken circuits of matroids (which can be express in terms of intersection numbers), we were motivated to consider an input polynomial with coefficients coming from intersections of divsiors on a variety. So, we considered the case where the input polynomial is the specialization $V(u)=(Au+B)^{d}$ of the volume polynomial associated to a pair of divsiors $A,B$ on a smooth projective variety. In general, the gamma vector has the same sign as $(-1)^{r}$ when the intersections $A^{k}B^{d-k}$ alternate in sign. Suppose that these intersections are nonnegative. When we are working with $\mathbb{P}^{n}$, we can relate the sign of the gamma vector to the relative “size” of $A$ compared to $B$ (since these correspond to integers) (Example 3.4. When the ratio is very large, the gamma vector is negative. If it is very small, it is alternating in sign like the negative ratio case. Positivity properties of divisors are also related to the bounds in a more general setting since the $A^{k}B^{d-k}$ form a log concave sequence when $A$ and $B$ are nef. This can be used to reduce upper and lower bound conditions related to alternating sums of positive multiples of such terms to those on the first or last pair of elements in a sequence. Note that the connections to upper/lower bounds studied in the previous section carry over to volume polynomials with some additional interpretations and simplifications. Finally, we give an explanation of the $\mathbb{P}^{n}$ case in terms of Segre classes of vector bundles and make use of these classes to connect the rational functions giving components of the gamma vector in the derivative expression to positivity questions related to Segre classes (Example 3.7).
Acknowledgments
I’d like to thank Lutz Warnke for pointing out a more recent reference [27] describing an elementary approach to the Lagrange inversion formula. This work was supported by Horizon Europe ERC Grant number: 101045750 / Project acronym: HodgeGeoComb.
@@ 1.1 An explicit formula involving Catalan numbers, binomial coefficients, and derivatives
(Lagrange inversion formula, Theorem 5.4.2 on p. 38 of [25] and Theorem 2.1.1 on p. 214 of [17], Theorem 1 on p. 944 of [27])
@@ 1.2 Local-global differences
Note that $r\leq\lfloor\frac{n}{2}\rfloor$. This means that $12r\leq n<n+$ 1 and $J^{\prime}(u)(u+1)^{2r}$ has a positive power of $u+1$ in the denominator and $(1+u)h^{\prime}(u)-nh(u)$ in the numerator. In the numerator, the coefficient of $v^{k}$ is $(k+1)h_{k+1}-(n-k)h_{k}$. Before starting more explicit computations, we make a heuristic comment on the interpreation of the derivative. In the case of $f$-polynomials, the derivative of the $f$-polynomial gives the sum of the $f$-polynomials of the links over the vertices of the given simplicial complex. This is a special case of the result below.
Definition 1.8.
(Definition 5.3.4 on p. 232 of [9], p. 60 of [24])
Let $\Delta$ be a simplicial complex and $F$ be a a subset of the vertex set of $\Delta$. The star of $\mathbf{F}$ is the set $\operatorname{st}_{\Delta}(F)=\{G\in\Delta:F\cup G\in\Delta\}$, and the link of $\mathbf{F}$ is the set $\operatorname{lk}_{\Delta}(F)=\{G:F\cup G\in\Delta,F\cap G=\emptyset\}$. In other words, we are removing $F$ from elements of $\operatorname{st}_{\Delta}(F)$.
@@ 2.1 Bounds involving translations
$\frac{a_{k+2r-1}}{a_{k}}\geq\frac{k\binom{2r-k-1}{r-1}}{(k+2r-1)\binom{k+r-1}{r-1}}$
@@
Since $B(t)=R(t+1)$, we have that
@@
$\Longleftrightarrow\frac{a_{k+2r-1}}{a_{k}}$ $\geq\frac{k\binom{2r-k-1}{r-1}}{(k+2r-1)\binom{k+r-1}{r-1}}$
in order to have $\gamma_{r}\leq 0$ for even $r$ as well. This is a sort of lower bound condition. ∎
@@ 2.2 Upper/lower bounds and general gamma vectors
The coefficient of $b_{0}$ inside the sum is $-d\binom{d-r-1}{r-1}$. If $k\geq 1$, the coefficient of $b_{k}$ inside is the sum of the index $i=k-1$ and $i=k$ terms of the sum given by
@@
$=(-1)^{k-1}\left(k\left(\binom{d-r-k}{r-k}-\binom{d-r-k-1}{r-k-1}\right)+d\binom{d-r-k-1}{r-k-1}\right)$
@@
where we used the identity $\binom{p}{q}-\binom{p-1}{q-1}=\binom{p-1}{q}$ in the last line. This means that
@@
      if $0\leq r\leq\frac{d}{3}$ or $r=\frac{d}{2}$. For $r=\frac{d}{2}$, we have that $\operatorname{sgn}(\gamma_{\frac{d}{2}})=(-1)^{r}$. This also holds for $\frac{d}{3}<r<\frac{d}{2}$ if
@@
      for all $10\leq k\leq r-$ 1, then $\gamma_{r}\leq 0$. If $r=\frac{d}{2}$, we have that $\gamma_{\frac{d}{2}}\leq 0$.
@@
Some examples where Part 1 would apply is for include ±$\pm\overline{\chi}_{M}(t)$ and a translate ±$\pm\overline{\chi}_{M}(t+1)$ for reduced characteristic polynomials of matroids $M$ (with sign depending on the parity of the rank of $M$). As for Part 2b), we can think of face vectors where the entries increase quickly (Theorem 2.9). Examples for individual indices coming from independence complexes of matroids are in equality cases of upper bounds on face vectors of independence complexes of matroids in work of Castillo–Samper (Theorem 3.3 on p. 10 of [11]). We can move in between cases 2a) and 2b) by looking at reciprocal polynomials.
@@
Given the simplification above, it is clear that Part 1 holds. Also, Lemma 2.5 directly implies the statements in Part 2 for $r=\frac{d}{2}$ since $(-1)^{r}r\gamma_{r}$ is a positive multiple of the alternating sum of the $b_{i}$ in this case. For the remaining indices in Part 2, we start by noting that the binomial terms in each sum of Proposition 2.4 give a decreasing sequence as the index increases since
@@
Finally, we consider Part 2b). Note that the coefficients of the $b_{k}$ are decreasing in the second sum in the first line of Proposition 2.4. In order for the terms to be increasing, we need to have
@@
If $0\leq r\leq\frac{d}{3}$, the coefficients of the first sum in the first line of Proposition 2.4 are also decreasing. So, the condition for the terms of the first sum to increase is to have
@@ 2.3 Gamma vectors and (modified) $h$-vector representations
We will focus on the expression from Proposition 2.4 since want to look at sums of positive terms and the older formula may involve some negative terms. For the second sum in Proposition 2.4, it suffices to have $b_{k}=f_{k-1}(\Delta)$ for some simplicial complex $\Delta$. It remains to consider the first sum. The main issue is that have a sum $11\leq k\leq r-$ 1 (meaning that 1 or 2 terms are missing) and $Q=r$ instead of $r-1$. To “complete” this to something coming from a simplicial complex, we can either write $\binom{d-r-k-1}{r-k}=\binom{d-r-k-1}{r-k-1}\cdot\frac{d-2r-1}{r-k}$ and add one additional term for $k=0$ or keep the given binomial coefficient $\binom{d-r-k-1}{r-k}$ and add 2 additional terms for $k=0$ and $k=r$. We can combine these observations with the following result:
@@ 3 Intersection numbers and volume polynomials
Also, $r\gamma_{r}=h(P)$ for some auxiliary $(d-r-2)$-dimensional simplicial poset $P$ when $d$ is sufficiently large and the $A^{k}B^{d-k}$ are increasing for $10\leq k\leq r-$ 1.
Remark 3.3.
The second part was motivated by $f$-vectors of independence complexes of matroids (use shellability and Proposition 3 on p. 3931 of [12], [1]) and $h$-vectors of broken circuit complexes of matroids and the analogue of this setup for divisors of the conormal Chow ring of a matroid (see [19], [11], [5], [4]).. Some related further discussion is in Part 2 of Remark 3.5.
@@
The second statement is an application of Theorem 2.9 as in Proposition 2.10. ∎
@@
[…]
@@
$r\gamma_{r}\text{``='' }[u^{r-1}]\frac{dB^{d-1}(A-B)}{(1+u)^{d-2r+1}(1-\rho u)^{d-1}}.$